OpenStudy (anonymous):

calcute the volume of a cyclinder in terms of x. with radius 2x and height 60-x.

OpenStudy (lyrae):

r = 2x h = 60-x Volume of a cylinder: V = pi * r^2 * h Substitute r and h with their respective values and simplify.